The Holtzman Corporation has assets of $387,000, current liabilities of $55,000, and long-term liabilities of $123,000. There is $37,500 in preferred stock outstanding; 20,000 shares of common stock have been issued. a. Compute book value (net worth) per share. (Round your answer to 2 decimal places.) Book value per share$ 8.58 b. If there is $26,600 in earnings available to common stockholders, and Holtzmans stock has a P/E of 17 times earnings per share, what is the current price of the stock? (Do not round intermediate calculations. Round your final answer to 2 decimal places.) Current price This is a numeric cell, so please enter numbers only. c. What is the ratio of market value per share to book value per share? (Do not round intermediate calculations. Round your final answer to 2 decimal places.) Market value to book value20.55 times
Answer #1

a)Book value per share = (Total shareholders equity - preferred equity)/no. Of outstanding shares

Total shareholders equity = total assets - total liabilities = total assets -( current liabilities+ long term liabilities)

=387000-(55000+123000) = $209000

Book value per share = (209000 - 37500)/20000 = $8.575  

b) Earnings available to common shareholders = $26000

Earnings per share = Earnings available to common shareholders/ no. Of outstanding shares

= 26600/20000 =$ 1.33

P/E =17×

Current Price per share = 17 × Earnings per share = 17× 1.33 =$ 22.61

c)market value per share = current price per share =$22.61

Ratio of market value per share to book value per share = $22.61/$8.575 = 2.6367 times